OpenStudy (anonymous):

Use the following to determine the distance A (13,7) B (-9,4).

OpenStudy (angela210793):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

D = sqrt ( ( A's x - B's x)^2 + (A's y - B's y ) ) D = sqrt ( ( 13 - (-9) )^2 + ( 7 - 4 )^2 ) D = sqrt ( ( 22)^2 + 3^2 ) D = sqrt ( ( 484 + 9 ) ) D = sqrt (493) D = 22,20

OpenStudy (angela210793):

All u have to do now is replace 13=x1 -9=x2 7=y1 and 4=y2